OpenStudy (anonymous):

A box contains twenty cards numbered 1 through 20. One card is to be randomly drawn from the box. What is the probability that the card drawn will show a multiple of 6, a multiple of 7, or a multiple of 8?

OpenStudy (anonymous):

By multiples, does it mean like 6,12,24 and so on and such? cause if so, it'd be 6 out of 20 cards that could be picked. Cause 6 + 6= 12, 7+7=14, and 8+8=16. You see what I mean? @Christiana90

OpenStudy (anonymous):

So what is the answer?

OpenStudy (anonymous):

what answers does it give you?

OpenStudy (anonymous):

I'd say 6/20

OpenStudy (anonymous):

1/20,1/3,7/20,1/2

OpenStudy (anonymous):

So there is not 6/20 on there

OpenStudy (anonymous):

Dang, I guess try 1/3 but I don't wanna cause u to get it wrong

OpenStudy (anonymous):

thank you

OpenStudy (accessdenied):

Here are all the cards we would be noting: [1] [2] [3] [4] [5] [6] [7] [8] [9] [10] [11] [12] [13] [14] [15] [16] [17] [19] [19] [20] The multiples of 6: [6], [12], and [18] The multiples of 7: [7] and [14] The multiples of 8: [8] and [16] Count your quantity for each: 3 multiples of 6, 2 of 7, and 2 of 8. So total: 3 + 2 + 2 = 7. Take the number 7 out of the total number of cards for your answer!